Please, help me with these:
8.)If a triangle has sides of lengths 5, 12, and 13, is the triangle right?
Yes
No
Sometimes
Not enough information
9.) If a triangle has sides of lengths 8, 10 and 13, is the triangle right?
Yes
No
Sometimes
Not enough information

Answer by ankor@dixie-net.com(22740)

8.)If a triangle has sides of lengths 5, 12, and 13, is the triangle right?
Does 5^2 + 12^2 = 13^2? If it does, it is a right triangle
Yes
:
9.) If a triangle has sides of lengths 8, 10 and 13, is the triangle right?
Does 8^2 + 10^2 = 13^2??
No